The radius of the earth's very nearly circular orbit around the sun is 1.50×1011 m.

Find the magnitude of the earth's centripetal acceleration as it travels around the sun. Assume a year of 365 days.

Express your answer to three significant figures and with appropriate units.
Answer #1

v = 2π*R/T

= 2π*1.5x10^11m/(365d*24h/d*3600s/h) = 3.0x10^4m/s

ω = v/R

= 3.0x10^4/1.5x10^11 = 2.0x10^-7 rad/s

a = r*ω^2

= 1.5x10^11m*(2.0x10^-7)^2 = 6.0x10^-3m/s^2